Simplify the following expressions:
(a) (cos π/4 + i sin π/4)(cos 3π/4 + i sin 3π/4), (b) (cos π/4 + i sin π/4)2
(cos π/6 + i sin π/6)
1
Expert's answer
2020-05-24T20:41:08-0400

Answers:

(a) "(cos\frac{\pi}{4}+i\sin\frac{\pi}{4})(cos\frac{3\pi}{4}+i\sin\frac{3\pi}{4})= (\frac{\sqrt{2}}{2}+i\frac{\sqrt{2}}{2})(-\frac{\sqrt{2}}{2}+i\frac{\sqrt{2}}{2})="

"=-1"

(b)"2(cos\frac{\pi}{4}+i\sin\frac{\pi}{4})=\sqrt{2}+i\sqrt{2}"

"cos\frac{\pi}{6}+i\,sin\frac{\pi}{6}=\frac{\sqrt{3}}{2}+i\frac{1}{2}"